OpenStudy (anonymous):

Simplify 2 times the 4th root of 80 .

OpenStudy (anonymous):



OpenStudy (compassionate):

Okay, so that "4" is saying, "what numbers, times itself four times, will equal 80." Once you find that it's pretty much simply multiplication.

OpenStudy (anonymous):

this stuff just confusing

OpenStudy (compassionate):

What number, times itself four times, equals 80.

OpenStudy (anonymous):

20?

OpenStudy (anonymous):

@Compassionate

OpenStudy (compassionate):

No. 20^4 =/= 80. You're finding a nth Root, in this case, the fourth root. 20 * 20 * 20 * 20 = 160,000

OpenStudy (compassionate):

It's kinda a trial and error thing. Keep plugging numbers into the 4th root and see if you get eighty. e..g, 4^4 = 4 * 4 * 4 * 4 = 256

OpenStudy (anonymous):

okay do you wanna see my answer choices

OpenStudy (compassionate):

Go ahead

OpenStudy (anonymous):